OpenStudy (anonymous):

Choose the correct simplification of the expression (xyz2)4. x5y5z6 x5y5z8 xyz16 x4y4z8

OpenStudy (anonymous):

D

OpenStudy (diyadiya):

\[\large (x~y~z^2)^4 = (x)^4(y)^4(z^2)^4 = x^4y^4 z ^{2\times 4} = x^4 y^4 z^8\] Is this clear?